A projectile with a mass of 20 g is fired from a gun at a 55 degree angle from the horizontal with a velocity of 550 m/s. If air resistance is neglected, what is its kinetic energy when it hits the ground?
Answer #1

when the projectile reaches the ground all of its energy is converted into kinetic energy so

K.E=1/2\times m\times v^{2}

=1/2\times (20/1000)\times 550^{2}=3025joule